find equation of the line if it passes through (2,3) and with x-intercept twice the y-intercept
Answer by stanbon(75887)

find equation of the line if it passes through (2,3) and with x-intercept twice the y-intercept
------------
x-intercept: (2k,0)
y-intercept: (0,k)
--------
slope = (k-0)/(0-2k) = -1/2
-----
Form: y = mx + b
Solve for "b":
3 = (-1/2)2 + b
b = 3+1 = 4
=======
Equation:
y = (-1/2)x + 4
==============================
Cheers,
Stan H.